OpenStudy (anonymous):

Help Please!! How do I calculate the kinetic friction coefficient on inclined plane? For example in this problem: The mass on the rough inclined plane of Length 3.5 meters reaches a velocity of 5 m/s at the bottom of the incline. calculate: a) potential energy at the top b) kinetic energy at the bottom c) work done by friction d) average friction force acting on the mass e) the coefficient of friction between the mass and the plane I know the answers to a and b.. it's the friction part I';m confused with.. Pls help me! :(

OpenStudy (anonymous):



OpenStudy (anonymous):

c) the work the friction does is the potential the top minus the kinetic at the bottom - ideally this would be zero, but the difference between the two is the work done by the friction (labeled as WNC -non conservative work) \[ E_i = E_f\] \[U_i = K_f+W_{NC}\] \[ W_{NC} = U_i-K_f\] d) \[W = Fd \cos \varphi\] where phi is the angle between the force of friction and the distance over which it acts (ph = 90). Use the work from c, and solve for F e) \[F=\mu_k N\] N is the normal force that you can find with a free body diagram. Use the force from d, and solve for mu ^_^

OpenStudy (anonymous):

thank you!